OpenStudy (amberlykhan):

medal will be given! simple 6th grade math! plz help!

OpenStudy (amberlykhan):

One part of green pigment mixed with 3 parts of yellow pigment creates lime green. Suppose the crayon factory is making lime-green crayons. They are going to use 140 total gallons of pigment. How many gallons of yellow and how many gallons of green will be used?

OpenStudy (mathmale):

I'd suggest you think of "3 parts of yellow and 1 part of green" as a ratio. Letting x be the number of gallons of green pigment used and y the number of gallons of yellow, \[\frac{ x }{ y }=\frac{ 1 }{ 3 }.\] You want 140 gallons total, right? Then x+y must = 140. Can you solve this set of simultaneous equations? Perhaps by using substitution?

OpenStudy (amberlykhan):

Yes, thank you.

OpenStudy (mathmale):

Good going! Cool.